lowing information for the next five questions: fferent shoe companies are competing over who can make the least manufacturing errors 1 takes a sample of 1250 shoes and finds that 23 of them have defects while Company 2 mple of 1500 shoes and finds that 63 of them have defects.
20. What is the pooled estimate for the proportion? a. 00184 b. 0042 c. 0.3127 e. None af the above 21 A new sample was taken, and a 99% confidence interval for the difference in the 2 i proportions. Which of the following could be It was found that the results were not statistically significant. the correct confidence interval? a. (-0.0032,1.5432) b. (0.0405, 0.8980) c (0.2138, 0.4325) d. (0.6789-0.2342) e. (2.4534, 3.4542) 22. A significance is run for the hypotheses: His μί-u2-0 and Ha: μ-μ2-0 was run. A 95% confidence interval was also constructed based on the same sample as the one above and the results were generated: (1.432,3.345). What does this mean for the significance test? a. The results are statistically significant- fail to reject the null hypothesis. b. The results are statistically significant- reject the null hypothesis. c. The results are practically and statistically significant - reject the null hypothesis. d. The results not are statistically significant- reject the null hypothesis. e The results not are statistically significant-fail to reject the null hypothesis.
Answer #1

20) \widehat p1 = 23/1250 = 0.0184

      \widehat p2 = 63/1500 = 0.042

The pooled proportion(P) = (\widehat p1 * n1 + \widehat p2 * n2)/(n1 + n2)

                                         = (0.0184 * 1250 + 0.042 * 1500)/(1250 + 1500)

                                         = 0.03127

Option - e) None of the above.

21) Option - A) (-0.0032, 1.5432)

Since the confidence interval contains 0, so we should not reject the null hypothesis.

so we can conclude that the results are not statistically significant.

22) Since the confidence interval does not contain 0, so we should reject the null hypothesis.

Option - b) The results are statistically significant - reject the null hypothesis.